# Verification and Synthesis of Compatible Control Lyapunov and Control Barrier Functions

### v1

@@ Abstract
Safety and stability are essential properties of control systems. Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) have been proposed to ensure safety and stability respectively. However, previous approaches typically verify and synthesize the CBFs and CLFs separately, satisfying their respective constraints, without proving that the CBFs and CLFs are compatible with each other, namely at every state, there exists control actions that satisfy both the CBF and CLF constraints simultaneously. There exists some recent works that synthesized compatible CLF and CBF, but relying on nominal polynomial or rational controllers, which is just a sufficient but not necessary condition for compatibility. In this work, we investigate verification and synthesis of compatible CBF and CLF independent from any nominal controllers. We derive exact necessary and sufficient conditions for compatibility, and further formulate Sum-Of-Squares program for the compatibility verification. Based on our verification framework, we also design an alternating nominal-controller-free synthesis method. We evaluate our method in a linear toy, a non-linear toy, and a power converter example.
@@ I Introduction
Safety and stability are essential for control systems that must be ensured to prevent catastrophic economic damage and loss of human life, while still achieve the desired goal [1, 2, 3, 4]. Lyapunov stability guarantees that all trajectories starting within the system’s region-of-attraction (RoA) will converge to the goal state. Meanwhile, safety properties are normally formulated as the positive invariance [5] of given regions in the state space.
@@
The importance of safety has motivated numerous methods such as artificial potential fields [6], reachability analysis[7], Hamilton–Jacobi analysis[8], model predictive control [9], and control barrier functions [10] to synthesize safe control strategies. Among these methods, energy-based methodologies such as barrier certificates [11] and Control Barrier Functions (CBFs) [10] have been proposed that construct an energy function that is positive if the system is safe, and then demonstrate safety by proving that the energy function remains positive for all time. Similarly, Lyapunov stability is certified by a Lyapunov function [12] or a Control Lyapunov function (CLF) [13], which remains non-negative and decreasing along any trajectories, and achieves the minimal value at the goal state. One major benefit of CLFs/CBFs is that they can be incorporated as constraints within optimization-based controllers, resulting in CLF-CBF control policies for joint safety and stability [5, 14]. Such techniques have been widely used in safety critical scenarios such as adaptive cruise control[5, 15], bipedal robot control [10, 16], and multi-waypoint navigation [3], with many different approaches to synthesize the CLF-CBF functions [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
A key challenge in CLF-CBF-based safe control policies is the compatibility of the CLF and CBF constraints, namely the existence of at least one admissible control action satisfying both CLF and CBF constraints simultaneously at every state in the safe region. Most previous approaches address incompatibility by relaxing the CLF constraints [5], and thus prioritizing safety over stability. However, this may sacrifice the stability and result in the system remaining in undesired states [32]. Recent efforts [33, 34] have utilized Sum-Of-Squares (SOS) methods [35, 36] to verify and synthesize compatible CLF and CBF based on nominal controllers. In such methods, a nominal controller is parameterized as a polynomial (or rational) function of state; these methods search for the CLF/CBF together with this nominal controller, such that this nominal controller simultaneously satisfies both the CLF and CBF constraints. The disadvantages of this approach is that it is only a sufficient but not necessary condition for compatibility, as there might exist a non-polynomial (or non-rational) compatible controller, which cannot be captured by the explicit parameterization. A related but orthogonal direction of research synthesizes closed-form controllers for CBF and CLF that are known to be compatible [37].
@@
In this paper, we study the problem of exact verification and synthesis of compatible CLF/CBF for all states in the safe region. We consider a nonlinear control-affine system, both with and without input limits, and derive necessary and sufficient conditions for compatibility using Farkas’ Lemma and the algebraic-geometric Postivstellensatz. We formulate the verification as a SOS program and further simplify it using S-procedure. To synthesize compatible CLF/CBF, we propose a bi-linear alternating method based on the simplified verification program. To summarize, this paper makes the following contributions.
@@
  • We derived the necessary and sufficient condition for CLF-CBF compatibility, and formulated a SOS program independent from any nominal controllers for compatibility verification.
@@
  • We next proposed an alternating method based on verification program to synthesis compatible CBF and CLF, also independent from any nominal controllers.
  • We demonstrated our method in a linear toy, a non-linear toy and a power converter examples.
@@ II Mathematical background
for some polynomials $p_{i}(\mathbf{x})$. A sos polynomial is non-negative. It can be verified that a polynomial is sos through convex optimization.
@@
$\mathbb{I}[\psi_{1},\ldots,\psi_{r}]=\left\{\sum_{l=1}^{r}{a_{i}(\mathbf{x})\psi_{l}(\mathbf{x})}:a_{1},\ldots,a_{r}\mbox{ polynomials}\right\}.$
@@
In Positivstellensatz, (2) is a necessary and sufficient condition for (1). On the other hand, (2) requires computing the cone $\Sigma$, which needs exponential number of polynomials $\alpha_{S}(\mathbf{x})$ (as $S$ permutes over all subsets of $\{1,\ldots,s\}$). To remedy this computation challenge, people often resort to S-procedure[36], which replaces the cone $\Sigma$ with the quadratic module $\mathcal{Q}$
@@
Eq. (2) in theorem 1 is a sufficient condition for (1) if we replace the cone $\theta\in\Sigma$ with the quadratic module $\theta\in\mathcal{Q}[\phi_{1},\ldots,\phi_{s}]$. [^1] Under some stronger assumptions on $\phi(\mathbf{x}),\psi(\mathbf{x})$, such as Archimedean, S-procedure also gives a necessary and sufficient condition [38].
@@ III-A System Model
where $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{n_{x}}$ is the state and $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{n_{u}}$ is the control input. We also assume that $f:\mathcal{X}\rightarrow\mathbb{R}^{n_{x}}$ and $g:\mathcal{X}\rightarrow\mathbb{R}^{n_{x}\times n_{u}}$ are polynomial functions of $\mathbf{x}$. The set of admissible control $\mathcal{U}$ is represented as a polyhedron $A\mathbf{u}\leq\mathbf{c},A\in\mathbb{R}^{p\times n_{u}},\mathbf{c}\in\mathbb{R}^{p}$. Note that this parameterization of $\mathcal{U}$ includes the entire set $\mathcal{R}^{n_{u}}$ (unconstrained control) as a special case, where $A$ and $\mathbf{c}$ are empty, hence we can model systems with or without input limits. A control policy $\mu:\mathcal{X}\rightarrow\mathcal{U}$ maps a state $\mathbf{x}$ to a control input $\mathbf{u}$. Without loss of generality we assume the equilibrium state is $\mathbf{x}=\mathbf{0}$. The state space $\mathcal{X}$ is divided into a safe region $\mathcal{X}_{s}$ and an unsafe region $\mathcal{X}_{u}=\mathcal{X}-\mathcal{X}_{s}$. Our goal is to design a control policy $\mu$ which renders $\mathcal{X}_{s}$ positively invariant, and the trajectories converging to $\mathbf{x}=\mathbf{0}$.
@@ III-B Control Barrier Function (CBF) and Control Lyapunov Function (CLF)
We consider safety constraints that are specified as super-level sets of differentiable functions $b:\mathcal{X}\rightarrow\mathbb{R}$. In particular, we consider sets $\mathcal{C}$ that satisfy
$\mathcal{C}$ =$=\{\mathbf{x}\;|\;b(\mathbf{x})\geq 0\}.$
If $\mathcal{C}\subseteq\mathcal{X}_{s}$ and $\mathcal{C}$ is positively invariant, then any trajectory with $\mathbf{x}(0)\in\mathcal{C}$ will be safe for all time $t$.
@@
A function $b(\mathbf{x})$ is a CBF if
$\sup_{\mathbf{u}\in\mathcal{U}}\{L_{f}b(\mathbf{x})+L_{g}b(\mathbf{x})\mathbf{u}+\kappa_{b}(b(\mathbf{x}))\}\geq 0$ (2)
for all $\mathbf{x}\in\mathcal{C}$, where $L_{f}b(\mathbf{x})=\frac{\partial b(\mathbf{x})}{\partial\mathbf{x}}\mathbf{f}(\mathbf{x})$, $L_{g}b(x)=\frac{\partial b(\mathbf{x})}{\partial\mathbf{x}}g(\mathbf{x})$, and $\kappa_{b}(\cdot)$ is an extended class $\kappa$ function [^2] Recall function $\kappa(.)$ is extended class $\kappa$ if it is monotone, increasing, and satisfies $\kappa(0)=0$..
A special class of Definition 1 consists of Exponential Control Barrier Functions, for which $\kappa_{b}(b(\mathbf{x}))=\kappa_{b}b(\mathbf{x})$ where $\kappa_{b}$ is a positive constant. Constraint (2) ensures the set $\mathcal{C}$ positive invariant for system represented by (1) [10].
@@
for all $\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$, where $\kappa_{V}(.)$ is also an extended class $\kappa$ function. This proves that the sub-level set $\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$ is an inner approximation of the region-of-attraction (ROA).
If a system satisfies constraint (3) for all time $t$, then the system can be stabilized to the origin [40]. Also, in this paper, we use exponential stability so that $\kappa_{V}(V(\mathbf{x}))=\kappa_{V}V(\mathbf{x})$, where $\kappa_{V}$ is a positive constant.
@@ IV-A Verification of compatibility
For a candidate CBF $b(\mathbf{x})$ and a candidate CLF $V(\mathbf{x})$, define 0-super-level set $\mathcal{C}=\{\mathbf{x}\;|\;b(\mathbf{x})\geq 0\}$ and the 1-sub-level set $\mathcal{D}:=\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$ (this 1-sub-level set will be an inner approximation of the region-of-attraction). We formally state the compatible CLF-CBF as follows
@@
Given positive constants $\kappa_{V}$, $\kappa_{b}$, a positive definite function $V(\mathbf{x})$, and a function $b(\mathbf{x})$ such that $\mathcal{C}\subseteq\mathcal{X}_{s}$, verify that $\forall\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$, there exists $\mathbf{u}$ such that
$\begin{bmatrix}L_{f}V(\mathbf{x})\\ -L_{f}b(\mathbf{x})\\ \mathbf{0}\end{bmatrix}+\begin{bmatrix}L_{g}V(\mathbf{x})\\ -L_{g}b(\mathbf{x})\\ A\end{bmatrix}\mathbf{u}\leq\begin{bmatrix}-\kappa_{V}V(\mathbf{x})\\ \kappa_{b}b(\mathbf{x})\\ \mathbf{c}\end{bmatrix}$ (4)
namely there exists $\mathbf{u}\in\mathcal{U}$ satisfying both the CBF condition (1) and the CLF condition (2) simultaneously. The region $\mathcal{C}\cap\mathcal{D}$ is certified to be safe and stabilizable.
We call the functions $V(\mathbf{x})$ and $b(\mathbf{x})$ compatible CLF and CBF if they satisfy the conditions of Problem 1, with $\mathcal{C}\cap\mathcal{D}$ as the compatible region.
IV-A1 Exact Verification

### v2

@@ Abstract
Safety and stability are essential properties of control systems. Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) are powerful tools to ensure safety and stability respectively. However, previous approaches typically verify and synthesize the CBFs and CLFs separately, satisfying their respective constraints, without proving that the CBFs and CLFs are compatible with each other, namely at every state, there exists control actions within the input limits that satisfy both the CBF and CLF constraints simultaneously. Ignoring the compatibility criteria might cause the CLF-CBF-QP controller to fail at runtime. There exists some recent works that synthesized compatible CLF and CBF, but relying on nominal polynomial or rational controllers, which is just a sufficient but not necessary condition for compatibility. In this work, we investigate verification and synthesis of compatible CBF and CLF independent from any nominal controllers. We derive exact necessary and sufficient conditions for compatibility, and further formulate Sum-Of-Squares programs for the compatibility verification. Based on our verification framework, we also design a nominal-controller-free synthesis method, which can effectively expands the compatible region, in which the system is guaranteed to be both safe and stable. We evaluate our method on a non-linear toy problem, and also a 3D quadrotor to demonstrate its scalability. The code is open-sourced at https://github.com/hongkai-dai/compatible_clf_cbf.
@@ I Introduction
Ensuring safety and stability are essential for control systems to prevent catastrophic economic damage and loss of human life, while still achieving the desired goal [1, 2, 3, 4]. Lyapunov stability guarantees that all trajectories starting within the system’s region-of-attraction (RoA) will converge to the goal state. Meanwhile, safety properties are normally formulated as the positive invariance [5] of given regions in the state space.
@@
The importance of safety has motivated numerous methods such as artificial potential fields [6], reachability analysis[7], Hamilton–Jacobi analysis[8], model predictive control [9], and control barrier functions [10] to synthesize safe control strategies. Among these methods, energy-based methodologies such as barrier certificates [11] and Control Barrier Functions (CBFs) [10] have been proposed that construct an energy function that is positive if the system is safe, and then demonstrate safety by proving that the energy function remains positive for all time. Similarly, Lyapunov stability is certified by a Lyapunov function [12] or a Control Lyapunov function (CLF) [13], which remains non-negative and decreasing along any trajectories, and achieves the minimal value at the goal state. One major benefit of CLFs/CBFs is that they can be incorporated as constraints within optimization-based controllers, resulting in CLF-CBF control policies for joint safety and stability [5, 14]. Such techniques have been widely used in safety critical scenarios such as adaptive cruise control[5, 15], bipedal robot control [10, 16], and multi-waypoint navigation [3], with many different approaches to synthesize the CLF and CBF functions [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
Figure 1: With the CLF-CBF-QP controller, the quadrotor converges to the desired goal (red sphere in the center) from different initial states, while avoiding the grey ground. The compatible CLF/CBF are synthesized by our algorithm. We also plot each trajectory of the quadrotor body origin with a different color.
@@
A key challenge in CLF-CBF-based safe control policies is the compatibility of the CLF and CBF constraints, namely the existence of at least one admissible control action satisfying both CLF and CBF constraints simultaneously at every state in the safe and stabilizable region. Most previous approaches address incompatibility by relaxing the CLF constraints [5], and thus prioritizing safety over stability. However, this may sacrifice the stability and result in the system remaining in undesired states [32]. Recent efforts [33, 34, 35] have utilized Sum-Of-Squares (SOS) methods [36, 37] to verify and synthesize compatible CLF and CBFs based on nominal controllers, for continuous-time control-affine systems with polynomial dynamics. In such methods, a nominal controller is parameterized as a polynomial (or rational) function of state; these methods search for the CLF/CBF together with this nominal controller, such that this nominal controller simultaneously satisfies both the CLF and CBF constraints. The disadvantage of this approach is that it is only a sufficient but not necessary condition for compatibility, as there might exist a non-polynomial (or non-rational) compatible controller, which cannot be captured by the explicit parameterization. A related but orthogonal direction of research synthesizes closed-form controllers for CBF and CLF that are known to be compatible [38].
@@
In this paper, we study the problem of exact verification and synthesis of compatible CLF/CBF for all states in the safe region. We consider a nonlinear control-affine system, both with and without input limits, and derive necessary and sufficient conditions for compatibility using Farkas’ Lemma and the algebraic-geometric Positivstellensatz. We formulate the exact verification as a SOS program and further simplify it using S-procedure. To synthesize compatible CLF/CBF, we propose a bi-linear alternating method based on the simplified verification program. To summarize, this paper makes the following contributions:
@@
  • We derive the necessary and sufficient condition for CLF-CBF compatibility, and formulate a SOS program independent from any nominal controllers for compatibility verification.
@@
  • We next propose an alternating method based on verification program to synthesis compatible CBF and CLF, also independent from any nominal controllers.
  • We demonstrate our method on a toy system and showcase its scalability on a 13-state quadrotor.
@@ II Mathematical background
for some polynomials $p_{i}(\mathbf{x})$. A sos polynomial is non-negative by definition. We can verify a polynomial being sos through convex optimization.
@@
$\mathbb{I}[\psi_{1},\ldots,\psi_{r}]\\ =\left\{\sum_{l=1}^{r}{a_{i}(\mathbf{x})\psi_{l}(\mathbf{x})}:a_{1},\ldots,a_{r}\mbox{ polynomials}\right\}.$
@@
In Positivstellensatz, (2) is a necessary and sufficient condition for (1). On the other hand, (2) requires computing the cone $\Sigma$, which needs exponential number of polynomials $\alpha_{S}(\mathbf{x})$ (as $S$ permutes over all subsets of $\{1,\ldots,s\}$). Besides, the “necessary” part is guaranteed only when the degrees of $\alpha_{S}(\mathbf{x})$ in the cone $\Sigma[\phi_{1},\ldots,\phi_{s}]$ and $a_{i}(\mathbf{x})$ in the ideal $\mathbb{I}[\psi_{1},\ldots,\psi_{r}]$ are sufficiently high, leading to excessively large optimization programs [39]. To remedy this computation challenge, people often resort to S-procedure[37], which replaces the cone $\Sigma$ with the quadratic module $\mathcal{Q}$
@@
Statement (2) in theorem 1 is a sufficient condition for (1) if we replace the cone $\theta\in\Sigma$ with the quadratic module $\theta\in\mathcal{Q}[\phi_{1},\ldots,\phi_{s}]$. [^1] Under some stronger assumptions on $\phi(\mathbf{x}),\psi(\mathbf{x})$, such as Archimedean, S-procedure also gives a necessary and sufficient condition [40].
To reduce the computation burden, in the S-procedure, we typically restrict both $\alpha_{i}(\mathbf{x})$ in the quadratic module, and $a_{i}(\mathbf{x})$ in the ideal to relatively low-degree polynomials.
@@ III-A System Model
where $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{n_{x}}$ is the state and $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{n_{u}}$ is the control input. We also assume that $f:\mathcal{X}\rightarrow\mathbb{R}^{n_{x}}$ and $g:\mathcal{X}\rightarrow\mathbb{R}^{n_{x}\times n_{u}}$ are polynomial functions of $\mathbf{x}$. The set of admissible control $\mathcal{U}$ is represented as a polyhedron $A\mathbf{u}\leq\mathbf{c},A\in\mathbb{R}^{p\times n_{u}},\mathbf{c}\in\mathbb{R}^{p}$. Note that this parameterization of $\mathcal{U}$ includes the entire set $\mathcal{R}^{n_{u}}$ (unconstrained control) as a special case, where $A$ and $\mathbf{c}$ are empty, hence we can model systems both with and without input limits. Without loss of generality we assume that the equilibrium state is $\mathbf{x}=\mathbf{0}$.
Many interesting dynamical systems don’t exhibit polynomial dynamics at first glance (namely $f(\mathbf{x}),g(\mathbf{x})$ are not polynomials). For example, robots with revolute joints typically use joint angle $\theta$ as the state, and its dynamics involves $\sin\theta$ and $\cos\theta$ functions, which are not polynomials. To overcome this, a widely-used technique is a change-of-variable trick: instead of using $\theta,\dot{\theta}$ as the state, we use $s=\sin\theta,c=\cos\theta$ and $\dot{\theta}$ as the state, with the time derivative as $\dot{s}=\dot{\theta}c,\dot{c}=-\dot{\theta}s$ both as polynomial functions of the new state $(s,c,\dot{\theta})$. Additionally we need to introduce the constraint $s^{2}+c^{2}=1$ on the state. As a result, we also include the algebraic equality constraint
$e(\mathbf{x})=0$ (1c)
in the system dynamics (where $e(\mathbf{x})=s^{2}+c^{2}-1$ for the revolute joint example). $e(\mathbf{x})$ is a polynomial function of $\mathbf{x}$. This technique on converting the non-polynomial dynamics to polynomial one has been used in previous SOS-based controller design[42, 43].
The state space $\mathcal{X}$ is divided into a safe region $\mathcal{X}_{s}$ and an unsafe region $\mathcal{X}_{u}=\mathcal{X}-\mathcal{X}_{s}$. Our goal is to formally certify that within the safe region $\mathcal{X}_{s}$, a control-invariant set such that all trajectories starting within this control-invariant set will converge to $\mathbf{x}=\mathbf{0}$.
@@ III-B Control Barrier Function (CBF) and Control Lyapunov Function (CLF)
We consider safety constraints that are specified as super-level sets of differentiable functions $h:\mathcal{X}\rightarrow\mathbb{R}$. In particular, we consider a set $\mathcal{C}$ parameterized as
$\mathcal{C}$ =$=\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}.$
If $\mathcal{C}\subseteq\mathcal{X}_{s}$ and $\mathcal{C}$ is control-invariant, then starting from any $\mathbf{x}(0)\in\mathcal{C}$, there exists control actions to keep the trajectory safe for all time $t\geq 0$.
@@
A function $h(\mathbf{x})$ is a CBF if
$\sup_{\mathbf{u}\in\mathcal{U}}\{L_{f}h(\mathbf{x})+L_{g}h(\mathbf{x})\mathbf{u}+\kappa_{h}(h(\mathbf{x}))\}\geq 0$ (2)
for all $\mathbf{x}\in\mathcal{C}$, where $L_{f}h(\mathbf{x})=\frac{\partial h(\mathbf{x})}{\partial\mathbf{x}}\mathbf{f}(\mathbf{x})$, $L_{g}h(\mathbf{x})=\frac{\partial h(\mathbf{x})}{\partial\mathbf{x}}g(\mathbf{x})$, and $\kappa_{h}(\cdot)$ is an extended class $\kappa$ function [^2] Recall function $\kappa(.)$ is extended class $\kappa$ if it is monotone, increasing, and satisfies $\kappa(0)=0$..
Constraint (2) ensures that the set $\mathcal{C}$ is control-invariant for systems represented by (1) [10].
@@
for all $\mathcal{D}=\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$, where $\kappa_{V}(.)$ is also an extended class $\kappa$ function. This proves that the sub-level set $\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$ is an inner approximation of the region-of-attraction (ROA).
To use our SOS tool, we assume that both the CLF $V(\mathbf{x})$ and CBF $h(\mathbf{x})$ are polynomial functions of $\mathbf{x}$. We also assume that $\kappa_{h}(\cdot),\kappa_{V}(\cdot)$ are polynomial. We will show that for a given pair of CBF/CLF candidates, we can certify the compatibility condition for general polynomial function $\kappa_{h}(\cdot),\kappa_{V}(\cdot)$. On the other hand, to synthesize the CBF/CLF, we consider a special class of Definition 1 and 2 consists of Exponential Control Barrier/Lyapunov Functions, for which $\kappa_{h}(h(\mathbf{x}))=\kappa_{h}h(\mathbf{x}),\kappa_{V}(V(\mathbf{x}))=\kappa_{V}V(\mathbf{x})$ where $\kappa_{h},\kappa_{V}$ are given positive constants.
@@ IV-A Verification of compatibility
For a candidate CBF $h(\mathbf{x})$ and a candidate CLF $V(\mathbf{x})$, define 0-super-level set $\mathcal{C}=\{\mathbf{x}\;|\;h(\mathbf{x})\geq 0\}$ and the 1-sub-level set $\mathcal{D}:=\{\mathbf{x}\;|\;V(\mathbf{x})\leq 1\}$. We formally state the compatible CLF-CBF as follows
@@
Given extended class $\kappa$ functions $\kappa_{V}(\cdot)$, $\kappa_{h}(\cdot)$, a positive definite function $V(\mathbf{x})$, and a function $h(\mathbf{x})$ such that $\mathcal{C}\subseteq\mathcal{X}_{s}$, verify that $\forall\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$, there exists $\mathbf{u}$ such that
$\begin{bmatrix}L_{f}V(\mathbf{x})\\ -L_{f}h(\mathbf{x})\\ \mathbf{0}\end{bmatrix}+\begin{bmatrix}L_{g}V(\mathbf{x})\\ -L_{g}h(\mathbf{x})\\ A\end{bmatrix}\mathbf{u}\leq\begin{bmatrix}-\kappa_{V}(V(\mathbf{x}))\\ \kappa_{h}(h(\mathbf{x}))\\ \mathbf{c}\end{bmatrix},$ (4)
namely there exists admissible $\mathbf{u}\in\mathcal{U}=\{\mathbf{u}|A\mathbf{u}\leq\mathbf{c}\}$ satisfying both the CBF condition (1) and the CLF condition (2) simultaneously. The region $\mathcal{C}\cap\mathcal{D}$ is control-invariant by definition, hence certified to be safe and stabilizable. We call the functions $V(\mathbf{x})$ and $h(\mathbf{x})$ compatible CLF and CBF if they satisfy the conditions of Problem 1, with $\mathcal{C}\cap\mathcal{D}$ as the compatible region.
IV-A1 Compatibility Verification